OpenStudy (anonymous):

need math help will give medal x - 7 = -19

OpenStudy (anonymous):

x = -12

OpenStudy (anonymous):

ty @kobe9115 thats wut i thought can u help me with more?

OpenStudy (anonymous):

1 = –(–w + 6)

OpenStudy (anonymous):

sry

OpenStudy (anonymous):

its 1= -(-w+6)

OpenStudy (anonymous):

w=7

OpenStudy (anonymous):

can u tell me why ?

OpenStudy (anonymous):

wouldnt it be -7?

OpenStudy (anonymous):

If you do 1= -(-w+6) then the w would be come positive, so it would be come 1=w-6, so 7-6=1.

OpenStudy (anonymous):

ready for the next one ?

OpenStudy (anonymous):

Sure, why not.

OpenStudy (anonymous):

(x+8) - 14 = -15

OpenStudy (anonymous):

It would become x-6=-15, which will give you x=-9

OpenStudy (anonymous):

yea thats wut i thought

OpenStudy (anonymous):

2 more

OpenStudy (anonymous):

@kobe9115

OpenStudy (anonymous):

erena parachutes from a plane at a height of 20,000 ft above sea level. Several minutes later she lands on the ground 15 ft below sea level . What was the total vertical distance that Serena parachuted?

OpenStudy (anonymous):

Wouldn't it be 20,000 feet? Since that's how for down she went?